OpenStudy (anonymous):

The sum of one third of a number and its reciprocal is the same as 49 divided by the number. Find the number. I believe the way it is set up is: 1/3x + 1/x = 49/x or x/3 + 1/x = 49/x

OpenStudy (anonymous):

okay that was wrong, you have \[\frac{x}{3}+\frac{1}{x}=\frac{49}{x}\] that is a good start

OpenStudy (anonymous):

multiply both sides by \(3x\) to get \[x^2+3=147\]

OpenStudy (anonymous):

this gives \(x^2=144\) and so \(x=12\) or \(x=-12\)

OpenStudy (anonymous):

Thanks again !!!!:D